# Self-Supervised Contrastive Learning is Approximately Supervised Contrastive Learning

### v1

@@ Appendix A Additional Experiments
Datasets.  We experiment with the following standard vision classification datasets - CIFAR10 and CIFAR100 [85], mini-ImageNet [86], and SVHN [87]. CIFAR10 and CIFAR100 both consist of 50000 training images and 10000 validation images with 10 classes and 100 classes, respectively, uniformly distributed across the dataset, i.e., CIFAR10 has 5000 samples per class and CIFAR100 has 500 samples per class. mini-ImageNet also has 5000 test images on top of 50000 train and 10000 validation images, with 100 of 1000 classes from ImageNet-1K [97] (at the original resolution). SVHN consists of digit classification data with 10 classes from real-world images, organized into "train" (73,257 samples), "test" (26,032 samples), and "extra" (531,131 samples) splits. For scalability verification, we combine the "train" and "extra" splits during training for validating Thm. 1 (shown in Fig. 2).
@@
Data augmentations.  We use the same augmentations as in SimCLR [5]. For experiments on mini-ImageNet and SVHN, we use the following pipeline: random resized cropping to $224\times 224$, random horizontal flipping, color jittering (brightness, contrast, saturation: $0.8$; hue: $0.2$), random grayscale conversion ($p=0.2$), and Gaussian blur (applied with probability $0.1$ using a $3\times 3$ kernel and $\sigma=1.5$). For CIFAR datasets, we adopt a similar pipeline with appropriately scaled parameters. The crop size is adjusted to $32\times 32$, and the color jitter parameters are scaled to saturation $0.4$, and hue $0.1$.
UMAP visualizations of clustering behaviors across datasets.  In Fig. 1, we show a visual comparison of the clusters formed by the learned representations of DCL and NSCL-based models during training for mini-ImageNet. In Fig. 6, we demonstrate that the behavior generalizes across different datasets, $i.e.$, CIFAR10, CIFAR100, and mini-ImageNet. We extract these embeddings using models trained for 2k epochs on their respective datasets.
A.1 Extension to MoCo
To verify the generality of our empirical findings in the main text, we repeat some primary experiments using the Momentum Contrast method, specifically MoCo v2 [6]. This section summarizes empirical results of MoCo, mirroring our analyses for SimCLR which we presented earlier.
MoCo setup.  We use the same architecture as with SimCLR. To train our model, we use SGD as an optimizer. We set momentum to $0.9$ and the weight decay to $1\mathrm{e}^{-4}$. All experiments are carried out with a batch size of $B=256$. The base learning rate is set to $0.03$ [27]. Similar to our SimCLR training strategy, we employ a warm-up phase [90] for the first 10 epochs, followed by a cosine learning rate schedule [91].
A.1.1 Results
Validating Thm. 1 during training.  We train MoCo using the DCL loss for 2k epochs, and evaluate both the DCL and NSCL losses on training and test datasets, along with the theoretical upper bound $\mathcal{L}^{\mathrm{NSCL}}(f)+\log(1+\tfrac{n_{max}\mathrm{e}^{2}}{N-n_{max}})$. We repeat the experiments illustrated in Fig. 2(top). In Fig. 7(top), we notice similar trends for MoCo as we have shown for SimCLR. The DCL loss consistently upper bounds the NSCL loss, with a narrowing gap with increase in number of classes $C$. The empirical $\mathcal{L}^{\mathrm{CL}}$ and $\mathcal{L}^{\mathrm{NSCL}}$ values remain tightly bounded by our proposed bound throughout training. We further validate that minimizing the DCL loss implicitly leads to low NSCL loss by repeating the experiments in Fig. 2(bottom). As shown in Fig. 7(bottom), $\mathcal{L}^{\mathrm{NSCL}}$ for both DCL and NSCL-trained models remain close. This confirms our earlier conclusion that optimizing DCL implicitly reduces the NSCL loss.
@@ A.1.1 Results
Comparing downstream performance.  We evaluate the quality of the learned representations from MoCo models via few-shot error analysis as earlier shown in Fig. 4. Specifically, we report the NCCC error on both train and test datasets. In Fig. 8(top) for MoCo, we evaluate NCCC error on the complete dataset as a function of number of shots per class ($m$) and show a comparison between DCL and NSCL-based models. In Fig. 8(bottom), we perform 2-way classification (averaged over 10 tasks) and compare the NCCC error with the bound from Cor. 1. We notice that the theoretical bound remains tight for increasing $m$. For example, at large $m$ (e.g., $m=10^{6}$) the test-time error bound drops down to $0.39$ for CIFAR100.
Similarly to our previous analysis, we also evaluate CDNV and directional CDNV in Fig. 9 for both DCL and NSCL-based MoCo models. We observe that while CDNV initially decreases for both models, the NSCL-trained models achieve substantially lower CDNV values as training progresses, whereas DCL-trained models maintain higher CDNV values. For directional CDNV, both the models achieve substantial reductions (an order of magnitude). For CIFAR10, directional CDNV for NSCL drops an additional order of magnitude compared to DCL.
Refer to caption Refer to caption
@@
Figure 8: (Top) $m$-shot error (Nearest Class-Centered Classifier (NCCC)) for $C$-way (all-class) classification. (Bottom) Two-way $m$-shot NCCC errors with the bound in Cor. 1. Same as Fig. 4 with MoCo training in place of SimCLR.
@@
Figure 9: CDNV and directional CDNV for both train and test data for DCL and NSCL-trained models over 2k epochs, for MoCo. Same as Fig. 5 with MoCo training in place of SimCLR.
@@ B.1 Batch-Based Contrastive Losses
In practice, contrastive learning objectives are implemented using batches of samples rather than full-dataset sums [5]. To formalize this, let $B\in\mathbb{N}$ denote a chosen batch size. We assume access to a dataset $S=\{(x_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times[C]$, where $x_{i}\in\mathcal{X}$ are inputs (e.g., images) and $y_{i}\in[C]$ are their class labels. For each sample $x_{i}$, let $x^{\prime}_{i}\sim\alpha(x_{i})$ be an independently generated augmentation from a distribution of augmentations $\alpha(x_{i})$. Given a batch =$\mathcal{B}=\{(x_{j_{t}},x^{\prime}_{j_{t}},y_{j_{t}})\}_{t=1}^{B}$ sampled with replacement from $S$, we define a per-sample contrastive loss for an anchor $(x_{i},y_{i})$ as:
@@
$-2(\log(2B)+2)\exp(-2B\epsilon^{2})~{}\leq~{}\mathcal{L}^{\mathrm{CL}}_{B}(f)-\mathcal{L}^{\mathrm{NSCL}}_{\bar{B}}(f)~{}\leq~{}\tfrac{\mathrm{e}^{2}(1+\epsilon C)}{C(1-\epsilon)-1}+2(\log(2B)+2)\exp(-2B\epsilon^{2}).$
@@
the corresponding embeddings. Next, let =$\mathcal{B}=\{(x_{j_{t}},x^{\prime}_{j_{t}},y_{j_{t}})\}_{t=1}^{B}$ be a random batch of $B$ samples (with replacement) from $S$ along with their augmentations. For the anchor $(x_{i},y_{i})$, define the per-sample contrastive loss by
@@
The proof is a technical modification of the proof of Prop. 7 in [82]. Fix a pair of classes $i\neq j$. Draw an independent dataset $\widehat{S}_{i}=\{x_{i,1},\dots,x_{i,m}\}\sim D_{i}^{m}$ and $\widehat{S}_{j}=\{x_{j,1},\dots,x_{j,m}\}\sim D_{j}^{m}$. Let $\widehat{\mu}_{i}=\tfrac{1}{m}\sum_{s=1}^{m}f(x_{i,s})$.
For a test point $x_{i}\sim D_{i}$ with embedding $z_{i}=f(x_{i})$, the NCC rule predicts class $j$ instead of $i$ if and only if $\lVert z_{i}-\widehat{\mu}_{j}\rVert_{2}\leq\lVert z_{i}-\widehat{\mu}_{i}\rVert_{2}$. By the triangle inequality and union bound:
$\Pr\left[\|z_{i}-\widehat{\mu}_{j}\|\leq\|z_{i}-\widehat{\mu}_{i}\|\right]$
$\leq~{}\Pr\left[\|z_{i}-\mu_{j}\|\leq\|z_{i}-\mu_{i}\|+\|\widehat{\mu}_{i}-\mu_{i}\|+\|\widehat{\mu}_{j}-\mu_{j}\|\right]$
$\leq~{}\Pr\left[\|z_{i}-\mu_{j}\|\leq\|z_{i}-\mu_{i}\|+2\Delta_{ij}~{}\lor~{}\|\widehat{\mu}_{i}-\mu_{i}\|\geq\Delta_{ij}~{}\lor~{}\|\widehat{\mu}_{j}-\mu_{j}\|\geq\Delta_{ij}\right]$
$\leq~{}\Pr\left[\|z_{i}-\mu_{j}\|\leq\|z_{i}-\mu_{i}\|+2\Delta_{ij}\right]+\Pr\left[\|\widehat{\mu}_{i}-\mu_{i}\|\geq\Delta_{ij}\right]+\Pr\left[\|\widehat{\mu}_{j}-\mu_{j}\|\geq\Delta_{ij}\right].$
By Markov’s inequality,
$\Pr\left[\|\widehat{\mu}_{i}-\mu_{i}\|\geq\Delta_{ij}\right]~{}\leq~{}\frac{\text{Var}(\widehat{\mu}_{i})}{\Delta_{ij}^{2}}~{}=~{}\frac{\text{Var}(z_{i})}{\Delta_{ij}^{2}m^{2}},$
and similarly
$\Pr\left[\|\widehat{\mu}_{j}-\mu_{j}\|\geq\Delta_{ij}\right]~{}\leq~{}\frac{\text{Var}(z_{j})}{\Delta_{ij}^{2}m^{2}}.$
Assume that $\|z_{i}-\mu_{j}\|\leq\|z_{i}-\mu_{i}\|+2\Delta_{ij}$ and $\|z_{i}-\mu_{i}\|\leq\alpha_{ij}$. Then,
$\|z_{i}-\mu_{i}\|^{2}+4\alpha_{ij}\Delta_{ij}+4\Delta_{ij}^{2}~{}$ $\geq~{}\|z_{i}-\mu_{i}\|^{2}+2\|z_{i}-\mu_{i}\|\cdot 2\Delta_{ij}+4\Delta_{ij}^{2}$
$=~{}(\|z_{i}-\mu_{i}\|+2\Delta_{ij})^{2}$
$\geq~{}\|z_{i}-\mu_{j}\|^{2}$
$=~{}\|(z_{i}-\mu_{i})-(\mu_{j}-\mu_{i})\|^{2}$
$=~{}\|z_{i}-\mu_{i}\|^{2}-2(z_{i}-\mu_{i})^{\top}(\mu_{j}-\mu_{i})+d_{ij}^{2}.$
Therefore, we have $2(z_{i}-\mu_{i})^{\top}(\mu_{j}-\mu_{i})~{}\geq~{}d_{ij}^{2}-4\alpha_{ij}\Delta_{ij}-4\Delta_{ij}^{2}$, which can also be written as follows
$(z_{i}-\mu_{i})^{\top}u_{ij}~{}\geq~{}\beta_{ij}~{}:=~{}\frac{1}{2}d_{ij}-\frac{2(\alpha_{ij}\Delta_{ij}+\Delta_{ij}^{2})}{d_{ij}}.$
Next we upper bound the terms $\Pr[(z_{i}-\mu_{i})^{\top}u_{ij}\geq\beta_{ij}]$ and $\Pr[\|z_{i}-\mu_{i}\|\geq\alpha_{ij}]$. Namely, by Markov’s inequality,
$\Pr[(z_{i}-\mu_{i})^{\top}u_{ij}\geq\beta_{ij}]~{}\leq~{}\frac{\text{Var}((z_{i}-\mu_{i})^{\top}u_{ij})}{\beta^{2}_{ij}},$
@@
$\Pr[\|z_{i}-\mu_{i}\|\geq\alpha_{ij}]~{}\leq~{}\frac{\text{Var}(z_{i})}{\alpha^{2}_{ij}}.$
Therefore, by combining everything together, we have:
$\Pr\left[\|z_{i}-\widehat{\mu}_{j}\|\leq\|z_{i}-\widehat{\mu}_{i}\|\right]~{}\leq~{}\frac{\text{Var}((z_{i}-\mu_{i})^{\top}u_{ij})}{\beta^{2}_{ij}}+\frac{\text{Var}(z_{i})}{\alpha^{2}_{ij}}+\frac{\text{Var}(z_{i})+\text{Var}(z_{j})}{\Delta_{ij}^{2}m^{2}}.$
We choose $\alpha_{ij}=\frac{d_{ij}\sqrt{m}}{\gamma_{1}}$ and $\Delta_{ij}=\frac{d_{ij}}{\gamma_{2}\sqrt{m}}$, then $\beta_{ij}=d_{ij}\left(\tfrac{1}{2}-\tfrac{2}{\gamma_{1}\gamma_{2}}-\tfrac{2}{\gamma^{2}_{2}m}\right)^{2}$. We have:
$\Pr\left[\|z_{i}-\widehat{\mu}_{j}\|\leq\|z_{i}-\widehat{\mu}_{i}\|\right]$
$\leq~{}\left(\tfrac{1}{2}-\tfrac{2}{\gamma_{1}\gamma_{2}}-\tfrac{2}{\gamma^{2}_{2}m}\right)^{-2}\frac{\text{Var}((z_{i}-\mu_{i})^{\top}u_{ij})}{d^{2}_{ij}}+\frac{\gamma^{2}_{1}\text{Var}(z_{i})}{md^{2}_{ij}}+\frac{\gamma^{2}_{2}(\text{Var}(z_{i})+\text{Var}(z_{j}))}{md^{2}_{ij}}$
$\leq~{}\left(\tfrac{1}{2}-\tfrac{2}{\gamma_{1}\gamma_{2}}-\tfrac{2}{\gamma^{2}_{2}m}\right)^{-2}\tilde{V}_{f}(D_{i},D_{j})+\frac{\gamma^{2}_{1}V_{f}(D_{j},D_{i})+2\gamma^{2}_{2}V_{f}(D_{i},D_{j})}{m}$
For the true class $i$ a mistake can occur in favor of any $j\neq i$. A union bound therefore gives
$\text{err}^{\mathrm{NCC}}_{m,D}(f){}$ $=~{}\frac{1}{C^{\prime}}\sum_{i=1}^{C^{\prime}}\Pr\bigl{[}\text{misclassify }x\sim D_{i}\bigr{]}$
$\leq~{}(C^{\prime}-1)\left[\left(\tfrac{1}{2}-\tfrac{2}{\gamma_{1}\gamma_{2}}-\tfrac{2}{\gamma^{2}_{2}m}\right)^{-2}\text{Avg}_{i\neq j}[\tilde{V}_{f}(D_{i},D_{j})]+\tfrac{\gamma^{2}_{1}+2\gamma^{2}_{2}}{m}\text{Avg}_{i\neq j}[V_{f}(D_{i},D_{j})]\right].$
Since the bound holds uniformly for all $\gamma_{1},\gamma_{2}>0$ such that $\tfrac{1}{2}-\tfrac{2}{\gamma_{1}\gamma_{2}}-\tfrac{2}{\gamma^{2}_{2}m}>0$, we can pick $\gamma_{1},\gamma_{2}$ such that their product is $a\geq 5$ and minimize $\gamma^{2}_{1}+2\gamma^{2}_{2}$. In this case, $\gamma_{1}=2^{1/4}\sqrt{a}$, $\gamma_{2}=2^{-1/4}\sqrt{a}$ and $\gamma^{2}_{1}+2\gamma^{2}_{2}=2\sqrt{2}a$. Let $\tilde{V}_{f}=\text{Avg}_{i\neq j}[\tilde{V}_{f}(D_{i},D_{j})]$ and $V_{f}=\text{Avg}_{i\neq j}[V_{f}(D_{i},D_{j})]$. The bound then becomes:
$\text{err}^{\mathrm{NCC}}_{m,D}(f){}$ $\leq~{}(C^{\prime}-1)\inf_{a\geq 5}\left[\left(\tfrac{1}{2}-\tfrac{2}{a}-\tfrac{2^{3/2}}{am}\right)^{-2}\tilde{V}_{f}+\tfrac{2\sqrt{2}a}{m}V_{f}\right].$ (8)
For $a=10$ we obtain that $\left(\tfrac{1}{2}-\tfrac{2}{a}-\tfrac{2^{3/2}}{am}\right)^{-2}<14$ and $28.52\sqrt{2}a<$ 28.5 which gives us the coefficients above.
Finally, $\mathrm{err}^{\mathrm{LP}}_{m,D}\leq\mathrm{err}^{\mathrm{NCC}}_{m,D}$ because the optimal linear probe cannot perform worse than the NCC classifier, since the NCC classifier is a special case of a linear probe. ∎

### v2

@@ Appendix A Additional Experiments
Datasets.  We experiment with the following standard vision classification datasets - CIFAR10 and CIFAR100 [krizhevsky2009learning], mini-ImageNet [NIPS2016_90e13578], Tiny-ImageNet [deep-learning-thu-2020], SVHN [37648] and ImageNet-1K [5206848]. CIFAR10 and CIFAR100 both consist of 50000 training images and 10000 validation images with 10 classes and 100 classes, respectively, uniformly distributed across the dataset, i.e., CIFAR10 has 5000 samples per class and CIFAR100 has 500 samples per class. mini-ImageNet also has 5000 test images on top of 50000 train and 10000 validation images, with 100 of 1000 classes from ImageNet-1K [5206848] (at the original resolution). Tiny-ImageNet contains 100000 images downsampled to $64\times 64$, with total 200 classes from IM-1K. Each class has 500 training, 50 validation, and 50 test images. SVHN consists of digit classification data with 10 classes from real-world images, organized into "train" (73,257 samples), "test" (26,032 samples), and "extra" (531,131 samples) splits. For scalability verification, we combine the "train" and "extra" splits during training for validating Thm. 1 (shown in Fig. 3). ImageNet-1K spans 1000 object classes and contains 1,281,167 training images, 50,000 validation images and 100,000 test images. We use "train" split for training and "validation" split for reporting test results.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(a) CIFAR10 (b) CIFAR100 (c) mini-ImageNet (d) Tiny-ImageNet
Figure 9: We reproduced Fig. 3 (top) for SimCLR with ViT-Base architecture. The loss-curves remain tightly bound as per Thm 1, independent of the encoder network architecture.
@@
Data augmentations.  We use the same augmentations as in SimCLR [pmlr-v119-chen20j]. For experiments on mini-ImageNet, and IM-1K, we use the following pipeline: random resized cropping to $224\times 224$, random horizontal flipping, color jittering (brightness, contrast, saturation: $0.8$; hue: $0.2$), random grayscale conversion ($p=0.2$), and Gaussian blur (applied with probability $0.1$ using a $3\times 3$ kernel and $\sigma=1.5$). For CIFAR datasets and SVHN, we adopt a similar pipeline with appropriately scaled parameters. The crop size is adjusted to $32\times 32$, and the color jitter parameters are scaled to saturation $0.4$, and hue $0.1$. For Tiny-ImageNet, we use the same saturation and hue values as CIFAR datasets with the crop size is adjusted to $64\times 64$.
UMAP visualizations of clustering behaviors.  In Fig. 1, we visualize the clustering behavior of DCL and NSCL-trained representations learned on mini-ImageNet using 2D UMAP [2018arXivUMAP] projections. We randomly selected 5 classes from mini-ImageNet and sampled 200 images per class, projecting their corresponding embeddings into 2D space using UMAP. As shown in Fig. 1, training with the NSCL loss results in tight and clearly separable clusters, as predicted in Thm. 2. Although DCL is label-agnostic, as shown in Thm. 1 and Fig. 3, DCL training implicitly minimizes the NSCL loss, which explains the visible clustering behavior seen in Fig. 1. In Fig. 8, we demonstrate that the behavior generalizes across different datasets, $i.e.$, CIFAR10, CIFAR100, and mini-ImageNet.
A.1 Extension to SimCLR (w/ ViT) and MoCo
To verify the generality of our empirical findings in the main text, we repeat some primary experiments with Vision Transformer (ViT) [dosovitskiy2021imageworth16x16words] in SimCLR, and also using the Momentum Contrast method, specifically MoCo v2 [chen2020improvedbaselinesmomentumcontrastive]. This section summarizes empirical results of SimCLR-ViT and MoCo, mirroring our analyses for SimCLR-ResNet-50 which we presented earlier.
SimCLR-ViT setup.  We adopt the ViT-Base (ViT-B/16) architecture, which consists of $12$ transformer layers, each with $12$ attention heads and a hidden dimension of $768$. For input images of size $224\times 224$, we use a patch size of $16\times 16$, resulting in $196$ tokens (plus a [CLS] token). The MLP hidden dimension is $3072$, and layer normalization is applied before each attention and MLP block. For CIFAR10 and CIFAR100, we follow the scaling procedure described in [yoshioka2024visiontransformers], adapting patch size and positional embeddings accordingly.
MoCo setup.  We use the same architecture as with SimCLR. To train our model, we use SGD as an optimizer. We set momentum to $0.9$ and the weight decay to $1\mathrm{e}^{-4}$. All experiments are carried out with a batch size of $B=256$. The base learning rate is set to $0.03$ [He_2020_CVPR]. Similar to our SimCLR training strategy, we employ a warm-up phase [goyal2017accurate] for the first 10 epochs, followed by a cosine learning rate schedule [loshchilov2016sgdr].
A.1.1 Results
@@
Figure 11: $m$-shot error (Nearest Class-Centered Classifier (NCCC)) for $C$-way (all-class) classification.
Refer to caption Refer to caption Refer to caption
(a) CIFAR10 (b) CIFAR100 (c) mini-ImageNet
Figure 12: Two-way $m$-shot NCCC and LP errors with ViT-Base to verify the bound in Cor. 1, as shown in Fig. 6 (bottom).
@@
Refer to caption Refer to caption Refer to caption
(c) CIFAR10 (d) CIFAR100 (e) mini-ImageNet
Figure 13: CDNV and directional CDNV analysis. Top Row: Results for a MoCo-trained ResNet-50 on CIFAR10 and CIFAR100. Bottom Row: Results for a SimCLR-trained ViT-Base on CIFAR10, CIFAR100, and mini-ImageNet.
Validating Thm. 1 during training.  We train SimCLR-ViT and MoCo using the DCL loss for 2k epochs, and evaluate both the DCL and NSCL losses on training and test datasets, along with the theoretical upper bound $\mathcal{L}^{\mathrm{NSCL}}(f)+\log(1+\tfrac{n_{max}\mathrm{e}^{2}}{N-n_{max}})$. We repeat the experiments illustrated in Fig. 3(top). In Fig. 9 and Fig. 10(top), we notice similar trends as we have shown for SimCLR. The DCL loss consistently upper bounds the NSCL loss, with a narrowing gap with increase in number of classes $C$. The empirical $\mathcal{L}^{\mathrm{CL}}$ and $\mathcal{L}^{\mathrm{NSCL}}$ values remain tightly bounded by our proposed bound throughout training. For MoCo, we further validate that minimizing the DCL loss implicitly leads to low NSCL loss by repeating the experiments in Fig. 3(bottom). As shown in Fig. 10(bottom), $\mathcal{L}^{\mathrm{NSCL}}$ for both DCL and NSCL-trained models remain close. This confirms our earlier conclusion that optimizing DCL implicitly reduces the NSCL loss.
Comparing downstream performance.  We evaluate the quality of the learned representations from MoCo models via few-shot error analysis as earlier shown in Fig. 6. Specifically, we report the NCCC error on both train and test datasets. In Fig. 11(top) for MoCo, we evaluate NCCC error on the complete dataset as a function of number of shots per class ($m$) and show a comparison between DCL and NSCL-based models. In Fig. 12 for SimCLR-ViT, we perform 2-way classification (averaged over 10 tasks) and compare the NCCC error with the bound from Cor. 1. We notice that the theoretical bound remains tight for increasing $m$.
Similarly to our previous analysis, we also evaluate CDNV and directional CDNV in Fig. 13 for both DCL and NSCL-based MoCo and SimCLR-ViT models. We observe that while CDNV initially decreases for both models, the NSCL-trained models achieve substantially lower CDNV values as training progresses, whereas DCL-trained models maintain higher CDNV values. For directional CDNV, both the models achieve substantial reductions (an order of magnitude). For CIFAR10, directional CDNV for NSCL drops an additional order of magnitude compared to DCL.
@@ B.1 Batch-Based Contrastive Losses
In practice, contrastive learning objectives are implemented using batches of samples rather than full-dataset sums [pmlr-v119-chen20j]. To formalize this, let $B\in\mathbb{N}$ denote a chosen batch size. We assume access to a dataset $S=\{(x_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times[C]$, where $x_{i}\in\mathcal{X}$ are inputs (e.g., images) and $y_{i}\in[C]$ are their class labels. For each sample $x_{i}$, let $x^{\prime}_{i}\sim\alpha(x_{i})$ be an independently generated augmentation from a distribution of augmentations $\alpha(x_{i})$. Given a batch $\mathcal{B}=\{(x_{j_{t}},x^{\prime}_{j_{t}},y_{j_{t}})\}_{t=1}^{B}$ sampled with replacement from $S$, we define a per-sample contrastive loss for an anchor $(x_{i},y_{i})$ as:
@@
$2(\log(2B)+2)\exp(-2B\epsilon^{2})~\leq~\mathcal{L}^{\mathrm{CL}}_{B}(f)-\mathcal{L}^{\mathrm{NSCL}}_{\bar{B}}(f)~\leq~\tfrac{\mathrm{e}^{2}(1+\epsilon C)}{C(1-\epsilon)-1}+2(\log(2B)+2)\exp(-2B\epsilon^{2}).$
@@
the corresponding embeddings. Next, let $\mathcal{B}=\{(x_{j_{t}},x^{\prime}_{j_{t}},y_{j_{t}})\}_{t=1}^{B}$ be a random batch of $B$ samples (with replacement) from $S$ along with their augmentations. For the anchor $(x_{i},y_{i})$, define the per-sample contrastive loss by
@@
We adapt Prop. 7 of [galanti2023generalizationboundsfewshottransfer]. Assume $d_{ij}:=\|\mu_{j}-\mu_{i}\|_{2}>0$ for all $i\neq j$ so that $u_{ij}:=(\mu_{j}-\mu_{i})/d_{ij}$ is well defined. Fix $i\neq j$. Draw independent samples $\widehat{S}_{i}=\{x_{i,1},\dots,x_{i,m}\}\sim D_{i}^{m}$ and $\widehat{S}_{j}=\{x_{j,1},\dots,x_{j,m}\}\sim D_{j}^{m}$, independent also of a fresh $x_{i}\sim D_{i}$. Let $\widehat{\mu}_{c}=\tfrac{1}{m}\sum_{s=1}^{m}f(x_{c,s})$ and $\mu_{c}=\mathbb{E}[f(x)\mid x\sim D_{c}]$ for $c\in\{i,j\}$, and write $z_{i}=f(x_{i})$. Set $\delta_{c}:=\widehat{\mu}_{c}-\mu_{c}$. For $c\in\{i,j\}$ define $z_{c}:=f(x_{c})$ with $x_{c}\sim D_{c}$, let $\Sigma_{c}:=\mathrm{Cov}(z_{c})$ and $v_{c}:=\operatorname{tr}(\Sigma_{c})$, and put
$\tilde{V}_{f}(D_{i},D_{j}):=\frac{u_{ij}^{\top}\Sigma_{i}u_{ij}}{d_{ij}^{2}},\qquad V_{f}(D_{i},D_{j}):=\frac{v_{i}+v_{j}}{d_{ij}^{2}}.$
Let $D:=\mu_{j}-\mu_{i}$, $d:=\|D\|_{2}$, and $A:=z_{i}-\tfrac{\mu_{i}+\mu_{j}}{2}$. A direct expansion gives
$\|z_{i}-\widehat{\mu}_{j}\|^{2}-\|z_{i}-\widehat{\mu}_{i}\|^{2}=d^{2}-2dX\;-\;2A^{\top}(\delta_{j}-\delta_{i})\;+\;D^{\top}(\delta_{j}+\delta_{i})\;+\;\big(\|\delta_{j}\|^{2}-\|\delta_{i}\|^{2}\big),$
where $X:=(z_{i}-\mu_{i})^{\top}u_{ij}$. Hence NCC predicts $j$ iff
$2X\;\geq\;d\;-\;\frac{2}{d}\,T\;+\;\frac{1}{d}\,S\;+\;\frac{1}{d}\,Q,\quad T:=A^{\top}(\delta_{j}-\delta_{i}),\;S:=D^{\top}(\delta_{j}+\delta_{i}),\;Q:=\|\delta_{j}\|^{2}-\|\delta_{i}\|^{2}.$ (8)
Fix $a\geq 5$ and $m\geq 6$, and set $\tau(a,m):=\tfrac{1}{2}-\tfrac{2}{a}-\tfrac{2^{3/2}}{am}\in(0,1/2)$, so $1-2\tau=\tfrac{4}{a}+\tfrac{2^{5/2}}{am}$. If
$\frac{2|T|}{d^{2}}+\frac{|S|}{d^{2}}+\frac{|Q|}{d^{2}}\ \leq\ 1-2\tau,$ (9)
then (8) implies $X\geq\tau d$, hence
$\{\mathrm{NCC\ predicts\ }j\}\ \subseteq\ \{X\geq\tau d\}\ \cup\ \Big\{\tfrac{2|T|}{d^{2}}+\tfrac{|S|}{d^{2}}+\tfrac{|Q|}{d^{2}}>1-2\tau\Big\}.$
Since $\mathbb{E}[X]=0$ and $\operatorname{Var}(X)=u_{ij}^{\top}\Sigma_{i}u_{ij}$,
$\Pr[X\geq\tau d]~\leq~\frac{\operatorname{Var}(X)}{\tau^{2}d^{2}}~=~\tau(a,m)^{-2}\,\tilde{V}_{f}(D_{i},D_{j}).$ (10)
By Markov and Cauchy–Schwarz,
$\Pr\Big[\tfrac{2|T|}{d^{2}}+\tfrac{|S|}{d^{2}}+\tfrac{|Q|}{d^{2}}>1-2\tau\Big]\ \leq\ \frac{2\,\mathbb{E}|T|+\,\mathbb{E}|S|+\,\mathbb{E}|Q|}{(1-2\tau)d^{2}}.$ (11)
Using independence of $(z_{i},\delta_{i},\delta_{j})$ and $\mathbb{E}[\delta_{c}]=0$,
$\mathbb{E}[T^{2}]=\frac{1}{m}\left(\operatorname{tr}\big(\Sigma_{i}(\Sigma_{i}+\Sigma_{j})\big)+\tfrac{d^{2}}{4}\,u_{ij}^{\top}(\Sigma_{i}+\Sigma_{j})u_{ij}\right),\qquad\mathbb{E}[S^{2}]=\frac{d^{2}}{m}\,u_{ij}^{\top}(\Sigma_{i}+\Sigma_{j})u_{ij}.$
Since $\operatorname{tr}(AB)\leq\operatorname{tr}(A)\operatorname{tr}(B)$ for PSD $A,B$ and $u^{\top}\Sigma u\leq\operatorname{tr}(\Sigma)$,
$\operatorname{tr}\big(\Sigma_{i}(\Sigma_{i}+\Sigma_{j})\big)\ \leq\ v_{i}\,(v_{i}+v_{j}),\qquad u_{ij}^{\top}(\Sigma_{i}+\Sigma_{j})u_{ij}\ \leq\ v_{i}+v_{j}.$
Hence, by Cauchy–Schwarz and $\sqrt{ab}\leq(a+b)/2$,
$\frac{2\,\mathbb{E}|T|}{d^{2}}$ $\leq\frac{2}{\sqrt{m}\,d^{2}}\sqrt{\,v_{i}(v_{i}+v_{j})+\frac{d^{2}}{4}(v_{i}+v_{j})}$
$\leq~\frac{1}{\sqrt{m}}\left(\sqrt{\frac{v_{i}+v_{j}}{d^{2}}}+2\,\frac{v_{i}+v_{j}}{d^{2}}\right)=\frac{1}{\sqrt{m}}\big(\sqrt{V_{ij}}+2V_{ij}\big),$
@@
$\frac{\mathbb{E}|S|}{d^{2}}~\leq~\frac{1}{\sqrt{m}}\sqrt{\frac{v_{i}+v_{j}}{d^{2}}}~=~\frac{1}{\sqrt{m}}\sqrt{V_{ij}},\qquad\mathbb{E}|Q|\ \leq\ \frac{v_{i}+v_{j}}{m}~=~\frac{V_{ij}\,d^{2}}{m}.$
Therefore
$\frac{2\,\mathbb{E}|T|+\,\mathbb{E}|S|+\,\mathbb{E}|Q|}{d^{2}}~\leq~\frac{1}{\sqrt{m}}\big(2\sqrt{V_{ij}}+2V_{ij}\big)~+~\frac{1}{m}V_{ij}.$ (12)
Combining the union bound, (10) and (12),
$\Pr(\mathrm{NCC\ predicts\ }j)~\leq~\tau(a,m)^{-2}\,\tilde{V}_{ij}\;+\;\frac{1}{1-2\tau(a,m)}\left(\frac{2}{\sqrt{m}}\sqrt{V_{ij}}+\frac{2}{\sqrt{m}}V_{ij}+\frac{1}{m}V_{ij}\right).$
Averaging over $i$ and summing over $j\neq i$ yields
$\mathrm{err}^{\mathrm{NCC}}_{m,D}(f)~\leq~(C^{\prime}-1)\left[\tau(a,m)^{-2}\,\tilde{V}_{f}\;+\;\frac{1}{1-2\tau(a,m)}\left(\frac{2}{\sqrt{m}}\,V^{s}_{f}+\frac{2}{\sqrt{m}}V_{f}+\frac{1}{m}V_{f}\right)\right],$
Since $(1-2\tau)^{-1}\leq a/4$, we obtain the simplified bound
$\mathrm{err}^{\mathrm{NCC}}_{m,D}(f)~\leq~(C^{\prime}-1)\inf_{a\geq 5}\left[\left(\tfrac{1}{2}-\tfrac{2}{a}-\tfrac{2^{3/2}}{am}\right)^{-2}\tilde{V}_{f}+\frac{a}{4}\left(\frac{2}{\sqrt{m}}V^{s}_{f}+\frac{2}{\sqrt{m}}V_{f}+\frac{1}{m}V_{f}\right)\right].$ (13)
By picking $a=16$ and using the assumption that $m\geq 10$ we get the desired bound. Finally, $\mathrm{err}^{\mathrm{LP}}_{m,D}\leq\mathrm{err}^{\mathrm{NCC}}_{m,D}$ since the Euclidean nearest-centroid rule is representable by a multiclass linear classifier. ∎